# Kai Shu, Lu Cheng, Huan Liu

### v3

@@ 4 Methodology
In this section, we present commonly adopted methods and tricks for LLM-as-a-judging, splitting them into tuning approaches (Section 4.1) and prompting strategies (Section 4.2).
@@ 4.1.1 Data Source
There are also some works that collect datasets for fine-grained judgment feedback.  Xu et al. (2023) introduce InstructScore, an explainable text generation evaluation metric, and curate the MetricInstruct dataset, which covers six text generation tasks and 23 datasets.  Liu et al. (2024a) collect ASPECTINSTRUCT, the first instruction-tuning dataset tailored for multi-aspect NLG evaluation, spanning 27 diverse evaluation aspects across 65 tasks.  Yue et al. (2023) first propose the evaluation of attribution and fine-tune judge LLMs with data from related tasks such as question answering, fact-checking, natural language inference, and summarization. One distinct approach is from Ke et al. (2024), which first prompts GPT-4 to produce the feedback and manually check its generated texts for each user query, revising them if necessary to improve the quality.
@@
While manually-labeled feedback is high-quality and accurately reflects human judgment preferences, it has limitations in terms of quantity and coverage. As a result, some researchers have turned to synthetic feedback as a data source for tuning judge LLMs. One approach in this direction relies on the judge LLMs themselves to generate the synthetic feedback. For example, Wu et al. (2024a) construct pairwise feedback for judgment enhancement by prompting the policy LLMs to evaluate their own judgments.  Wang et al. (2024f) prompt the LLM to generate a “noisy” version of the original instruction and use the corresponding response to this corrupted instruction as the inferior response.  Wang et al. (2024a) prompt GPT-4-Turbo to generate multiple pieces of evidence based on the original evidence for each instance, categorizing them into completely irrelevant evidence, partially irrelevant evidence and highly related evidence to train a hallucination judgment LLMs.  Park et al. (2024) build OFFSETBIAS, a pairwise preference dataset that leverages GPT-4 to generate bad, off-topic and erroneous responses and perform difficulty filtering. For safety judging, Xie et al. (2024a) adopt GPT-4 as the classifier to map each data point to a predefined safety category to train an automated evaluator. Different from previous works, Li et al. (2024e) adopt GPT-4 to synthesize both pairwise and pointwise data to train a generative judge LLM. For pointwise data, they adopt a “divide-and-conquer” strategy, where two critiques are collected from GPT-4 for a single response, combined into a more comprehensive critique, and a final rating is provided Following them, Kim et al. (2024) use GPT-4 to augment the preference learning dataset with detailed human evaluation criteria and verbal feedback. In the multi-modal domain, Xiong et al. (2024b) propose LLaVA-Critic and adopt GPT-4o to generate reasons behind given scores or preference judgments for training data construction. Moreover, By harnessing both explicit human instruction and the implicit knowledge of GPT-4, Xu et al. (2023) fine-tune a judge LLM based on LLaMA, producing both a score for generated text and a human-readable diagnostic report.  Zhu et al. (2023) introduce JudgeLM and propose a comprehensive, large-scale, high-quality dataset containing task seeds, LLMs-generated answers, and GPT-4-generated judgments for fine-tuning high-performance judges.
@@ 4.1.2 Tuning Techniques
Supervised fine-tuning (SFT) is the most commonly used approach to facilitate the judge LLMs to learn from pairwise Wang et al. (2024h); Li et al. (2024e); Wang et al. (2023b); Zhu et al. (2023) or pointwise Xie et al. (2024a); Wang et al. (2023b); Yue et al. (2023) judgment data. Among many works that adopted SFT, Vu et al. (2024) propose a supervised multitask training to tune their Foundational Large Autorater Models (FLAMe) across multiple mixed datasets of various tasks. To equip the judge LLM with both pairwise and pointwise judging capabilities, Kim et al. (2024) novelly propose joint training and weight merging approaches during the tuning stage and find the latter does not improve evaluation performances in the majority of cases. To obtain a judge model that can not only generate responses but also compare pairwise preferences, Lee et al. (2024) devise Judge-augmented Supervised Fine-tuning (JSFT) with an augmented preference learning dataset. During the training phase, Ke et al. (2024) enhance their model by adding simplified prompts to distinguish different parts of inputs and augment pairwise training data by swapping the order of two generated texts and exchanging the corresponding content in critiques.  Xu et al. (2023) further fine-tune their INSTRUCTSCORE model on self-generated outputs to optimize feedback scores, resulting in diagnostic reports that are better aligned with human judgment.  Liu et al. (2024a) also propose a two-stage supervised fine-tuning approach, first applying vanilla instruction tuning to equip the model with the ability to follow instructions for diverse evaluations. They then perform further tuning with auxiliary aspects to enrich the training process, incorporating an additional instruction-tuning stage to leverage potential connections to the target evaluation aspect.
@@
Preference learning is closely aligned with judgment and evaluation tasks, especially comparative and ranking judgment. In addition to works that directly adopt or augment preference learning datasets for supervised fine-tuning judge LLMs, several studies apply preference learning techniques to enhance LLMs’ judging capabilities. To enhance the quality of judgment provided by HALU-J, Wang et al. (2024a) further tune it with Directed Preference Optimization (DPO) Rafailov et al. (2023) after the SFT stage under the multiple-evidence setting. Similarly, Park et al. (2024) apply DPO with synthetic “bad” responses that contain critical errors but exhibit stylistic qualities favored by judge models, helping to mitigate bias in the judge LLMs.  Wu et al. (2024a) novelly propose meta-rewarding, which leverages the policy LLMs to judge the quality of their own judgment and produce pairwise signals for enhancing the LLMs’ judging capability. This concept is also adopted by Wang et al. (2024f), who propose self-taught evaluators that use corrupted instructions to generate suboptimal responses as inferior examples for preference learning. Recently, Hu et al. (2024) propose Themis, an LLM dedicated to NLG evaluation, which has been trained with designed multi-perspective consistency verification and rating-oriented preference alignment methods. Li et al. (2024n) propose PORTIA, an alignment-based approach designed to mimic human comparison behavior to calibrate position bias in an effective manner.
@@ 4.2.1 Swapping Operation
Previous studies have demonstrated that LLM-based judges are sensitive to the positions of candidates, and the quality ranking of candidate responses can be easily manipulated by merely altering their order in the context Wang et al. (2023c); Raina et al. (2024); Zhu et al. (2023). To mitigate this positional bias and establish a more fair LLM judging system, the swapping operation Zheng et al. (2023) has been introduced and widely adopted. This technique involves invoking the judge LLM twice, swapping the order of the two candidates in each instance. In evaluations, if the results are inconsistent after the swap, it is labeled a "tie," indicating that the LLM is unable to confidently distinguish the quality of the candidates Zheng et al. (2023). Several studies have also incorporated swapping operations in self-alignment Lee et al. (2023); Sun et al. (2024); Lee et al. (2024) to obtain more accurate pairwise feedback from the judge LLM. Zhu et al. (2024a) proposed a CoT-like prompting technique to mitigate the positional bias by asking the model to first provide all pairwise ranking, then summarize with a ranking list.
@@ 4.2.2 Rule Augmentation
Rule-augmented prompting involves embedding a set of principles, references, and evaluation rubrics directly within the prompt for the judge LLM. This approach is commonly employed in LLM-based evaluations, where judge LLMs are guided to assess specific aspects Li et al. (2024e); Bai et al. (2023a); Yu et al. (2024b); Qian et al. (2024) and provided with detailed rubrics Gao et al. (2023); Kim et al. ; Wang et al. (2024e); Murugadoss et al. (2024); Li et al. (2024k, h) to ensure a fair comparison. A distinct approach is seen in Liu et al. (2024b), where the judge LLM is prompted to generate its own scoring criteria through in-context learning on a set of few-shot examples. In alignment with LLM-as-a-judge, Bai et al. (2022) first propose to introduce a list of principles (e.g., helpfulness, harmlessness, honesty) for the judge LLM to compare the two candidates more precisely and directionally. Following them, subsequent works Lee et al. (2023, 2024); Guo et al. (2024); Sun et al. (2024); Beigi et al. (2024a) enhance this principle-driven prompting by incorporating more detailed explanations for each aspect of the principle or rubric. Besides, both Li and Qiu (2023) and Li et al. (2024c) propose to prompt LLMs to retrieve appropriate demonstrations/ knowledge triples based on the candidates’ helpfulness in solving specific problems. To obtain diverse responses from LLMs, Lahoti et al. (2023) prompt multiple LLMs to judge the diversity of each candidate and select the most diverse one for further polishing. Zhang et al. (2024f) propose RevisEval, which leverages the self-correction capabilities of LLMs to adaptively revise the response, then treat the revised text as the principles for the subsequent evaluation. Recently, Dong et al. (2024) investigate the reliability of LLM-as-a-personalized-judge, providing persona as a part of principles for LLMs to make personalized judgments.
@@ 4.2.3 Multi-agent Collaboration
Accessing results from a single LLM judge may not be reliable due to various biases inherent in LLMs Wang et al. (2023c); Liusie et al. (2024); Zhu et al. (2023); Koo et al. (2023a); Liu et al. (2023c). To address this limitation, Li et al. (2023) introduced the Peer Rank (PR) algorithm, which considers each peer LLM’s pairwise preferences for all answer pairs and produces a final ranking of models. Building on this foundation, several architectures and techniques for multi-agent LLMs have emerged, including mixture-of-agent Zhang et al. (2023b), role play Wu et al. (2023), debating Chan et al. (2023); Zhang et al. (2024e), and voting Zhu et al. (2024b). Jung et al. (2024) proposed Cascaded Selective Evaluation, where less expensive models serve as initial judges, escalating to stronger models only when necessary Beigi et al. (2024a). Additionally, some works apply multi-agent collaboration for alignment data synthesis, leveraging multiple LLM judges to refine responses Arif et al. (2024) or provide more accurate pairwise feedback Li et al. (2024i). Recnetly, Li et al. (2024l) propose the MATEval, where all agents are played by LLMs like GPT-4. The MATEval framework emulates human collaborative discussion methods, integrating multiple agents’ interactions to evaluate open-ended text.
@@ 4.2.5 Multi-turn Interaction
In evaluation, a single response may not provide enough information for an LLM judge to thoroughly and fairly assess each candidate’s performance. To address this limitation, multi-turn interactions are commonly adopted to offer a more comprehensive evaluation. Typically, the process begins with an initial query or topic, followed by dynamically interacting between the judge LLM and candidate models. Bai et al. (2023b) propose a multi-round setting where the evaluator takes on an interviewer role, posing increasingly sophisticated follow-up questions based on the models’ previous answers. Similarly, Yu et al. (2024b) introduce KIEval, a Knowledge-grounded Interactive Evaluation framework, which novelly incorporates an LLM-powered interactor to enable dynamic, contamination-resilient assessments. Additionally, some approaches facilitate debates among candidates in a multi-round format. For example, Zhao et al. (2024c) design a framework where two LLMs engage in a multi-round peer battle around a query, allowing their true performance differences to surface. Moniri et al. (2024) propose an automated benchmarking system where LLMs debate, with the final assessment carried out by another LLM judge.
@@ 5.1.1 Open-ended Generation Tasks
Open-ended generation refers to tasks where the generated content is expected to be safe, accurate, and contextually relevant, though there isn’t a single “correct” answer. Such tasks include dialog response generation, summarization, story generation, and creative writing Badshah and Sajjad (2024); Bai et al. (2023a); Kumar et al. (2024); Zeng et al. ; Yang et al. (2024). Unlike conventional metrics-based evaluation methods, LLM-as-a-judge provides a more nuanced, adaptable, and customized evaluation. As noted by Zheng et al. (2023), LLMs like GPT-4 perform comparably to humans when judging open-ended text generation. In practice, LLM-as-a-judge has been applied to evaluate outputs from a single model to compare outputs from multiple models in a competitive setting. For example, Gao et al. (2023) use ChatGPT to perform human-like summarization evaluation. Similarly, Wu et al. (2023) propose a comparison-based framework to let LLMs act as judges with multiple role-playing to evaluate the summarization quality in a specific dimension and generate its evaluations.
Modern LLMs are good at generating detailed and long-form responses. However, as the output length increases, so does the likelihood of hallucinations. To better understand this phenomenon, Cheng et al. (2023) and Zhang et al. (2024d) introduce an evaluation method that uses GPT-4 to judge whether generated outputs include logically structured yet nonsensical statements. Wang et al. (2024a) propose a critique-based judging system that evaluates hallucinations by selecting pertinent evidence and providing in-depth critiques. Beyond hallucinations, a significant concern is the generation of harmful (e.g., encouragement of self-suicide) and unsafe (e.g., guidance of illegal activity) responses by LLMs. Addressing this, Li et al. (2024g) introduce MD-Judge and MCQ-Judge for evaluating safety-related QA pairs, particularly focusing on queries crafted to provoke unsafe responses. This approach supports a seamless and reliable evaluation. However, an overly cautious stance toward unsafe queries can lead to excessive refusal responses, hindering normal functionality and negatively affecting user experience. To explore this issue, Xie et al. (2024a) conduct a meta-evaluation across various LLM-as-a-judge frameworks, assessing the refusal tendencies of current LLMs in response to potentially unsafe queries. Besides, Yu et al. (2024a) introduce an LLM-based answer extractor to accurately pinpoint critical part of answer from the open-ended generation. Additionally, An et al. (2023) propose L-Eval, a LLM-as-a-judge framework to institute a more standardized evaluation for long context language models.

### v4

@@ 4 Methodology
In this section, we present commonly adopted methods and tricks for LLM-as-a-judge, splitting them into tuning approaches (Section 4.1) and prompting strategies (Section 4.2).
@@ 4.1.1 Data Source
There are also some works that collect datasets for fine-grained judgment feedback Deshpande et al. (2024).  Xu et al. (2023) introduce InstructScore, an explainable text generation evaluation metric, and curate the MetricInstruct dataset, which covers six text generation tasks and 23 datasets.  Liu et al. (2024a) collect ASPECTINSTRUCT, the first instruction-tuning dataset tailored for multi-aspect NLG evaluation, spanning 27 diverse evaluation aspects across 65 tasks.  Yue et al. (2023) first propose the evaluation of attribution and fine-tune judge LLMs with data from related tasks such as question answering, fact-checking, natural language inference, and summarization. One distinct approach is from Ke et al. (2024), which first prompts GPT-4 to produce the feedback and manually check its generated texts for each user query, revising them if necessary to improve the quality.
@@
While manually-labeled feedback is high-quality and accurately reflects human judgment preferences, it has limitations in terms of quantity and coverage. As a result, some researchers have turned to synthetic feedback as a data source for tuning judge LLMs. One approach in this direction relies on the judge LLMs themselves to generate the synthetic feedback. For example, Wu et al. (2024a) construct pairwise feedback for judgment enhancement by prompting the policy LLMs to evaluate their own judgments.  Wang et al. (2024f) prompt the LLM to generate a “noisy” version of the original instruction and use the corresponding response to this corrupted instruction as the inferior response.  Wang et al. (2024a) prompt GPT-4-Turbo to generate multiple pieces of evidence based on the original evidence for each instance, categorizing them into completely irrelevant evidence, partially irrelevant evidence and highly related evidence to train a hallucination judgment LLMs.  Park et al. (2024) build OFFSETBIAS, a pairwise preference dataset that leverages GPT-4 to generate bad, off-topic and erroneous responses and perform difficulty filtering. For safety judging, Xie et al. (2024a) adopt GPT-4 as the classifier to map each data point to a predefined safety category to train an automated evaluator. Different from previous works, Li et al. (2024e) adopt GPT-4 to synthesize both pairwise and pointwise data to train a generative judge LLM. For pointwise data, they adopt a “divide-and-conquer” strategy, where two critiques are collected from GPT-4 for a single response, combined into a more comprehensive critique, and a final rating is provided Following them, Kim et al. (2024b) use GPT-4 to augment the preference learning dataset with detailed human evaluation criteria and verbal feedback. In the multi-modal domain, Xiong et al. (2024b) propose LLaVA-Critic and adopt GPT-4o to generate reasons behind given scores or preference judgments for training data construction. Moreover, By harnessing both explicit human instruction and the implicit knowledge of GPT-4, Xu et al. (2023) fine-tune a judge LLM based on LLaMA, producing both a score for generated text and a human-readable diagnostic report.  Zhu et al. (2023) introduce JudgeLM and propose a comprehensive, large-scale, high-quality dataset containing task seeds, LLMs-generated answers, and GPT-4-generated judgments for fine-tuning high-performance judges.
@@ 4.1.2 Tuning Techniques
Supervised fine-tuning (SFT) is the most commonly used approach to facilitate the judge LLMs to learn from pairwise Wang et al. (2024h); Li et al. (2024e); Wang et al. (2023b); Zhu et al. (2023) or pointwise Xie et al. (2024a); Wang et al. (2023b); Yue et al. (2023) judgment data. Among many works that adopted SFT, Vu et al. (2024) propose a supervised multitask training to tune their Foundational Large Autorater Models (FLAMe) across multiple mixed datasets of various tasks. To equip the judge LLM with both pairwise and pointwise judging capabilities, Kim et al. (2024b) novelly propose joint training and weight merging approaches during the tuning stage and find the latter does not improve evaluation performances in the majority of cases. To obtain a judge model that can not only generate responses but also compare pairwise preferences, Lee et al. (2024) devise Judge-augmented Supervised Fine-tuning (JSFT) with an augmented preference learning dataset. During the training phase, Ke et al. (2024) enhance their model by adding simplified prompts to distinguish different parts of inputs and augment pairwise training data by swapping the order of two generated texts and exchanging the corresponding content in critiques.  Xu et al. (2023) further fine-tune their INSTRUCTSCORE model on self-generated outputs to optimize feedback scores, resulting in diagnostic reports that are better aligned with human judgment.  Liu et al. (2024a) also propose a two-stage supervised fine-tuning approach, first applying vanilla instruction tuning to equip the model with the ability to follow instructions for diverse evaluations. They then perform further tuning with auxiliary aspects to enrich the training process, incorporating an additional instruction-tuning stage to leverage potential connections to the target evaluation aspect.
@@
Preference learning is closely aligned with judgment and evaluation tasks, especially comparative and ranking judgment. In addition to works that directly adopt or augment preference learning datasets for supervised fine-tuning judge LLMs, several studies apply preference learning techniques to enhance LLMs’ judging capabilities. To enhance the quality of judgment provided by HALU-J, Wang et al. (2024a) further tune it with Directed Preference Optimization (DPO) Rafailov et al. (2023) after the SFT stage under the multiple-evidence setting. Similarly, Park et al. (2024) apply DPO with synthetic “bad” responses that contain critical errors but exhibit stylistic qualities favored by judge models, helping to mitigate bias in the judge LLMs.  Wu et al. (2024a) novelly propose meta-rewarding, which leverages the policy LLMs to judge the quality of their own judgment and produce pairwise signals for enhancing the LLMs’ judging capability. This concept is also adopted by Wang et al. (2024f), who propose self-taught evaluators that use corrupted instructions to generate suboptimal responses as inferior examples for preference learning. Recently, Hu et al. (2024) propose Themis, an LLM dedicated to NLG evaluation, which has been trained with designed multi-perspective consistency verification and rating-oriented preference alignment methods. Li et al. (2024o) propose PORTIA, an alignment-based approach designed to mimic human comparison behavior to calibrate position bias in an effective manner.
@@ 4.2.1 Swapping Operation
Previous studies have demonstrated that LLM-based judges are sensitive to the positions of candidates, and the quality ranking of candidate responses can be easily manipulated by merely altering their order in the context Wang et al. (2023c); Raina et al. (2024); Zhu et al. (2023). To mitigate this positional bias and establish a more fair LLM judging system, the swapping operation Zheng et al. (2023) has been introduced and widely adopted. This technique involves invoking the judge LLM twice, swapping the order of the two candidates in each instance. In evaluations, if the results are inconsistent after the swap, it is labeled a "tie," indicating that the LLM is unable to confidently distinguish the quality of the candidates Zheng et al. (2023). Several studies have also incorporated swapping operations in self-alignment Lee et al. (2023); Sun et al. (2024a); Lee et al. (2024) to obtain more accurate pairwise feedback from the judge LLM. Zhu et al. (2024a) proposed a CoT-like prompting technique to mitigate the positional bias by asking the model to first provide all pairwise ranking, then summarize with a ranking list.
@@ 4.2.2 Rule Augmentation
Rule-augmented prompting involves embedding a set of principles, references, and evaluation rubrics directly within the prompt for the judge LLM. This approach is commonly employed in LLM-based evaluations, where judge LLMs are guided to assess specific aspects Li et al. (2024e); Bai et al. (2023a); Yu et al. (2024c); Qian et al. (2024) and provided with detailed rubrics Gao et al. (2023b); Kim et al. ; Wang et al. (2024e); Murugadoss et al. (2024); Li et al. (2024l, h) to ensure a fair comparison. A distinct approach is seen in Liu et al. (2024b), where the judge LLM is prompted to generate its own scoring criteria through in-context learning on a set of few-shot examples. In alignment with LLM-as-a-judge, Bai et al. (2022) first propose to introduce a list of principles (e.g., helpfulness, harmlessness, honesty) for the judge LLM to compare the two candidates more precisely and directionally. Following them, subsequent works Lee et al. (2023, 2024); Guo et al. (2024); Sun et al. (2024a); Beigi et al. (2024a) enhance this principle-driven prompting by incorporating more detailed explanations for each aspect of the principle or rubric. Besides, both Li and Qiu (2023) and Li et al. (2024c) propose to prompt LLMs to retrieve appropriate demonstrations/ knowledge triples based on the candidates’ helpfulness in solving specific problems. To obtain diverse responses from LLMs, Lahoti et al. (2023) prompt multiple LLMs to judge the diversity of each candidate and select the most diverse one for further polishing. Zhang et al. (2024g) propose RevisEval, which leverages the self-correction capabilities of LLMs to adaptively revise the response, then treat the revised text as the principles for the subsequent evaluation. Recently,some works investigate the reliability of LLM-as-a-personalized-judge Dong et al. (2024); Seraj et al. (2024); Kim and Baek (2024), providing persona as a part of principles for LLMs to make personalized judgments.
@@ 4.2.3 Multi-agent Collaboration
Accessing results from a single LLM judge may not be reliable due to various biases inherent in LLMs Wang et al. (2023c); Liusie et al. (2024); Zhu et al. (2023); Koo et al. (2023a); Liu et al. (2023c). To address this limitation, Li et al. (2023a) introduced the Peer Rank (PR) algorithm, which considers each peer LLM’s pairwise preferences for all answer pairs and produces a final ranking of models. Building on this foundation, several architectures and techniques for multi-agent LLMs have emerged, including mixture-of-agent Zhang et al. (2023b), role play Wu et al. (2023), debating Chan et al. (2023); Zhang et al. (2024f), and voting Zhu et al. (2024c); Rahmani et al. (2024). Jung et al. (2024) proposed Cascaded Selective Evaluation, where less expensive models serve as initial judges, escalating to stronger models only when necessary Beigi et al. (2024a). Additionally, some works apply multi-agent collaboration for alignment data synthesis, leveraging multiple LLM judges to refine responses Arif et al. (2024) or provide more accurate pairwise feedback Li et al. (2024i). Recnetly, Li et al. (2024m) propose the MATEval, where all agents are played by LLMs like GPT-4. The MATEval framework emulates human collaborative discussion methods, integrating multiple agents’ interactions to evaluate open-ended text.
@@ 4.2.5 Multi-turn Interaction
In evaluation, a single response may not provide enough information for an LLM judge to thoroughly and fairly assess each candidate’s performance. To address this limitation, multi-turn interactions are commonly adopted to offer a more comprehensive evaluation. Typically, the process begins with an initial query or topic, followed by dynamically interacting between the judge LLM and candidate models. Bai et al. (2023b) propose a multi-round setting where the evaluator takes on an interviewer role, posing increasingly sophisticated follow-up questions based on the models’ previous answers. Similarly, Yu et al. (2024c) introduce KIEval, a Knowledge-grounded Interactive Evaluation framework, which novelly incorporates an LLM-powered interactor to enable dynamic, contamination-resilient assessments. Additionally, some approaches facilitate debates among candidates in a multi-round format. For example, Zhao et al. (2024c) design a framework where two LLMs engage in a multi-round peer battle around a query, allowing their true performance differences to surface. Moniri et al. (2024) propose an automated benchmarking system where LLMs debate, with the final assessment carried out by another LLM judge.
@@ 5.1.1 Open-ended Generation Tasks
Open-ended generation refers to tasks where the generated content is expected to be safe, accurate, and contextually relevant, though there isn’t a single “correct” answer. Such tasks include dialog response generation, summarization, story generation, and creative writing Badshah and Sajjad (2024); Bai et al. (2023a); Kumar et al. (2024); Zeng et al. ; Yang et al. (2024). Unlike conventional metrics-based evaluation methods, LLM-as-a-judge provides a more nuanced, adaptable, and customized evaluation. As noted by Zheng et al. (2023), LLMs like GPT-4 perform comparably to humans when judging open-ended text generation. In practice, LLM-as-a-judge has been applied to evaluate outputs from a single model to compare outputs from multiple models in a competitive setting. For example, Gao et al. (2023b) use ChatGPT to perform human-like summarization evaluation. Similarly, Wu et al. (2023) propose a comparison-based framework to let LLMs act as judges with multiple role-playing to evaluate the summarization quality in a specific dimension and generate its evaluations.
Modern LLMs are good at generating detailed and long-form responses. However, as the output length increases, so does the likelihood of hallucinations. To better understand this phenomenon, Cheng et al. (2023) and Zhang et al. (2024d) introduce an evaluation method that uses GPT-4 to judge whether generated outputs include logically structured yet nonsensical statements. Wang et al. (2024a) propose a critique-based judging system that evaluates hallucinations by selecting pertinent evidence and providing in-depth critiques. Beyond hallucinations, a significant concern is the generation of harmful (e.g., encouragement of self-suicide) and unsafe (e.g., guidance of illegal activity) responses by LLMs. Addressing this, Li et al. (2024g) introduce MD-Judge and MCQ-Judge for evaluating safety-related QA pairs, particularly focusing on queries crafted to provoke unsafe responses. This approach supports a seamless and reliable evaluation. However, an overly cautious stance toward unsafe queries can lead to excessive refusal responses, hindering normal functionality and negatively affecting user experience. To explore this issue, Xie et al. (2024a) conduct a meta-evaluation across various LLM-as-a-judge frameworks, assessing the refusal tendencies of current LLMs in response to potentially unsafe queries. Besides, Yu et al. (2024a) introduce an LLM-based answer extractor to accurately pinpoint critical part of answer from the open-ended generation. Additionally, An et al. (2023) propose L-Eval, a LLM-as-a-judge framework to institute a more standardized evaluation for long context language models. This concept are followed by Bai et al. (2024) and they propose to leverage LLM-as-a-judge to filter evaluation data for long-context LLMs.